OpenStudy (anonymous):

Solve the following |x+8|≤5

OpenStudy (zzr0ck3r):

\(|a|\le b \iff -b\le a \le b\) so you have \(|x+8|\le 5 \implies -5\le x+8\le 5\) now subtract \(8\) from every term.

OpenStudy (anonymous):

so -13≤x≤-3. How do i graph that on a number line?

OpenStudy (anonymous):

How do I graph my answer on a number line?

OpenStudy (texaschic101):

| x + 8 | <= 5 (x + 8) <= 5 --- subtract 8 from both sides x <= 5 - 8 x <= -3 -(x + 8) <= 5 -x - 8 <= 5 --- add 8 to both sides -x <= 5 + 8 -x <= 13 -- multiply by -1 to make x positive, and change inequality sign x >= -13 so you have : -13 <= x <= -3

OpenStudy (texaschic101):

to graph on a number line... when there is equal signs, the dots are closed. less then (<) means shaded to the left and greater then (>) means shaded to the right. So your number line will look like this : closed circles on -3 and -13, shading in between.

OpenStudy (zzr0ck3r):

is there an echo in here?

OpenStudy (anonymous):